What volume (mL) of a concentrated solution of sodium hydroxide (6.00 M) must be diluted to 200.0 mL to make a 0.880 M solution of sodium hydroxide? Question 37 options: A) 50.0 B) 29.3 C) 176 D) 26.4 E) 2.64
Answer #1

we know the formula, M1V1 = M2V2

where M1 = initial concentration of the solution

V1 = initial volume of the solution

M2 = final concentration of the solution

V2 = final volume of the solution

given, V1 = to be determined

M1 = 6.00 M

V2 =200mL

M2 = 0.880M

from the formula we can write,

V1 = (M2V2)/ M1 = (200* 0.880) / 6 = 29.33 mL

therefore the initial volume of NaOH solution required is 29.33mL